HIGH-RESOLUTION MICROWAVE SPECTROSCOPY OF IMINOSILICON, HNSI

M. C. MCCARTHY, P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, U.S.A. and School of Engineering and Applied Sciences, Harvard University, 29 Oxford Street, Cambridge, MA 02138, U.S.A.; FILIPPO TAMASSIA, Dipartimento di Chimica Fisica ed Inorganica, Universitá di Bologna, V.le Risorgimento 4, I-40136 Bologna, Italy; \underline{S. THORWIRTH}, Max-Planck-Institut für Radioastronomie, 53121 Bonn, Germany, and I. Physikalisches Institut, Universität zu Köln, 50937 Köln, Germany.

By means of Fourier transform microwave spectroscopy of a supersonic beam, the fundamental rotational transition of isotopic and vibrationally-excited iminosilicon HNSi has been detected. In addition to seven isotopic species, vibrational satellite transitions from more than 30 vibrationally-excited states, including the three fundamental modes, have been detected. Those from v_{2} are particularly intense, enabling detection of transitions from as high as (0, 22^{0}, 0) (i.e. ∼10,000 cm^{−1} above ground). At high spectral resolution, well-resolved nitrogen quadrupole structure has been observed in nearly every transition. Excitation of v_{1} or v_{3} changes eQq(N) little, but eQq(N) systematically decreases with increasing excitation of the v_{2} bend, from a value of 0.376(5) MHz for (0, 0^{0}, 0) to -2.249(5) MHz for (0, 20^{0}, 0). With the large amount of new data in hand, it has been possible to determine more precise vibration-rotation constants and an improved semi-empirical structure for this triatomic molecule. An unsuccessful search for HSiN, a highly polar isomer calculated to lie nearly 3 eV above HNSi, is also reported.